OpenStudy (anonymous):

CAN SOMEONE HELP ME PLZZZZZZZZZZZZZ If a line crosses the y-axis at (0, -4) and has a slope of -2, what is the equation of the line? 2x − y = -4 2x + y = -4 2x + y = 4 2x − y = 4

OpenStudy (anonymous):

its going down so it would be 2x-y and then your y crosses at -4 =)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can u help me some more

OpenStudy (anonymous):

yeah =)

OpenStudy (anonymous):

A company that makes nails finds that it would sell 42 boxes of nails per month if they are priced at $20 per box. If the company lowers the price to $10 per box, then more people would purchase nails and the company could expect to sell 52 boxes per month. Find the equation of the line for the demand equation if x represents the number of boxes sold and y is the price of a box of nails. y = 62 + x y = 42 − x y = 42 + x y = 62 − x

OpenStudy (anonymous):

@MarziaJackson$

OpenStudy (anonymous):

nd the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0. slope = −14 and y-intercept = −94 slope = 14 and y-intercept = −94 slope = 14 and y-intercept = 94 slope = −14 and y-intercept = 94 Done

OpenStudy (anonymous):

Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0. slope = −14 and y-intercept = −94 slope = 14 and y-intercept = −94 slope = 14 and y-intercept = 94 slope = −14 and y-intercept = 94

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

if im cought on here im sol =)

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

k im back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0. slope = −14 and y-intercept = −94 slope = 14 and y-intercept = −94 slope = 14 and y-intercept = 94 slope = −14 and y-intercept = 94

OpenStudy (anonymous):

1st add 5 to both sides ...

OpenStudy (anonymous):

x- 4y +4

OpenStudy (anonymous):

= 5

OpenStudy (anonymous):

idk ..

OpenStudy (anonymous):

im lost =(

OpenStudy (anonymous):

i need to find the slope an the y intercept